If U = {0,1,2,3,4,5,6,7,8,9}, A = {3,6,9} and B = {4,8}, find:

a. A intersection B

b. A union B

c. A complement

d. B complement

e. (A union B)complement

1
Expert's answer
2022-06-07T12:39:37-0400

a.

There is no common elements for "A = \{3,6,9\}" and "B = \{4,8\}." Then


"A\cap B=\{\}"

b.

Take all elements from "A=\{3,6,9\}" and add elements "4, 8" from "B"


"A\cup B=\{3.4,6,8,9\}"

c.

All elements which are in "U = \{0,1,2,3,4,5,6,7,8,9\}," but are not in "A = \{3,6,9\}" we write in "A^C"


"A^C=\{0,1,2,4,5,7,8\}"

d. All elements which are in "U = \{0,1,2,3,4,5,6,7,8,9\}," but are not in "B= \{4,8\}" we write in "B^C"


"B^C=\{0,1,2,3,5,6,7,9\}"

e. All elements which are in "U = \{0,1,2,3,4,5,6,7,8,9\}," but are not in

"A\cup B=\{3.4,6,8,9\}" we write in "(A\cup B)^C"


"(A\cup B)^C=\{0,1,2,5,7\}"